OpenStudy (anonymous):

Please help :// Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x².

OpenStudy (anonymous):

i should make a template for this it is like the 6th time i have seen this exact problem

OpenStudy (gitface):

Must all be the same course ;)

OpenStudy (anonymous):

Haha really? could you help me tho? && yeah probably

OpenStudy (anonymous):

here is my example \[\large \frac{x^7\times x^3}{x^8}=\frac{x^{10}}{x^8}=x^{10-8}=x^2\] i am sure you can come up with others

OpenStudy (anonymous):

Eh Alright, Thank you :)

OpenStudy (anonymous):

clear enough? you can mimic it and find many more right?

OpenStudy (anonymous):

Yes :) It gave me an idea on how to do it, thank you very much!

OpenStudy (anonymous):

yw